Two capacitor of capacitance C1 and C2 connected in series are joined to two other capacitors of capacitance of capacitances C3 and C4, which are connected in parallel .determine the equivalent capacitance of the circuit .


1
Expert's answer
2022-02-04T07:44:39-0500

The two capacitors conected in series

"C_s=\frac{C_1C_2}{C_1+C_2}"



Two capacitors connected in parallel

"C_p= C_3+C_4"


"C_{total}=\frac{C_1+C_2}{C_1C_2}+\frac{1}{C_3+C_4}\\=\frac{C_1C_2+C_1C_3+C_1C_4+C_2C_3+C_2C_4}{(C_1C_2)(C_3+C_4)}"